I think this is the right topic. I need help with my Quadratic sequences homework, as I have a very bad teacher, and nearly no one in my class understands what he says.
I need help in finding the nth term for these sequences:
a)7,8,10,13,17
b)2,3,5,8,12
c)1,3,7,13,21
d)5,8,14,23,35
Please help, and if you could show your workings, that would help me a lot.

Answer by stanbon(75887)

I need help in finding the nth term for these sequences:
---
comment: In each case you need to "see" the pattern.
------------------------------------------------------------
a)7,8,10,13,17
a(1) = 7
a(2) = 7+1 = 8
a(3) = 8+2 = 10
a(4) = 10+3 = 13
a(5) = 13+4 = 17
....
a(n) = a(n-1)+(n-1) where a(1)=7 and n>=2
-------------------------------------------------
b)2,3,5,8,12
I think that 12 should be 5+8=13
Then a(n) = a(n-2)+a(n-1) where a(1)=2 and a(2)=3
-------------------------------------
c)1,3,7,13,21
a(n) = 1+2(n-1) for n = 0,1,2,3.....
------------------------------------
d)5,8,14,23,35
a(1) = 5
a(2) = a(1)+3
a(3) = a(2)+2(3)
a(4) = a(3)+3(3)
a(5) = a(4)+4(3)
---
a(n) = a(n-1)+3(n-1)
==========================
Cheers,
Stan H.